# Last Update: Steptember 24, 2025

### v1

@@
Last Update: Steptember 24, 2025
@@
Jerry Yao-Chieh Hu†‡∗111jhu@u.northwestern.edu; jhu@ensemblecore.ai  Xiwen Zhang§∗222xiz284@ucsd.edu  Weimin Wu333wwm@u.northwestern.edu  Han Liu†♯444hanliu@northwestern.edu
**footnotetext: These authors contributed equally to this work. Part of the work done during JH’s internship at Ensemble AI.
@@ 1 Introduction
Nevertheless, while Dao and Gu (2024) conjecture that analogues duality should hold for diagonal state-space models, there exist no formal treatment to the best of our knowledge. We give a concrete diagonal state-space duality, and provide the regarding computation algorithm.
@@
  • General Diagonal SSMs (Sections˜4.1 and 4.2). We extend SSD beyond the simple $(\text{scalar})\times I_{N}$ state matrix to general diagonal state matrices. This enlarges the class of SSMs under the duality (from a single exponential decay to $N$ separate diagonal dynamics), thereby supporting richer sequence dynamics.
@@
  • Higher-Rank Equivalence (Section˜A.1). We prove the conjecture of (Dao and Gu, 2024) that the equivalence between semiseparable matrices and matrices with sequential state-space representation holds not only for the rank-1 case, but also for general rank $N$. In other words, an SSM of state dimension $N$ corresponds to an $N$-semiseparable attention matrix.
@@
We provide related work in Section˜2, and background in Section˜3. We show our main theory in Section˜4, and the limitations of structured state-space duality in Section˜5.
@@ 2 Related Work
SSMs aim to model long-range dependencies with linear-time computation. S3 first introduced the idea of parameterizing state-space models with structured matrices to enable efficient sequence modeling (Gu et al., 2021a). S4 introduces a state-space layer with stable diagonalization and fast convolution, which enabled long-context training and inference (Gu et al., 2021b). S5 simplifies the design with a single multi-input multi-output SSM and preserved $O(T)$ scaling (Smith et al., 2023). Mamba adds input-dependent gating on the SSM projections and achieved strong accuracy with linear-time sequence modeling (Gu and Dao, 2023). These works establish SSMs as competitive sequence models and motivate analyses that compare their expressive power to attention.
@@ 4.3 Computational Complexity of Diagonal SSD
We define the functions $f:\mathbb{R}^{T}\times\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{T\times d}$ by $f(x,Y)_{:,s}=x\odot(Y_{:,s})$ for all $s\in[d]$,, and $g:\mathbb{R}^{T}\times\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{T\times d}$ by $g(x,Y)_{1,:}=Y_{1,:}$ and $g(x,Y)_{t+1,:}=x_{t+1}\cdot g(x,Y)_{t,:}+Y_{t+1,:}$ for $t\in[T-1]$. Consider the SSM layer with state dimension $N$ defined by (3.2), where each $A^{t}$ is a diagonal matrix. The recurrence relation (3.2) also has representation
@@
  $Z^{n}\leftarrow f(b^{n},X)\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $\mathcal{O}(NTd)$
  $H^{n}\leftarrow g(a^{n},Z^{n})\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $\mathcal{O}(NTd)$
  $Y^{n}\leftarrow f(c^{n},H^{n})\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $\mathcal{O}(NTd)$
  $Y\leftarrow\sum_{n=1}^{N}Y^{n}$ // Time $\mathcal{O}(NTd)$
@@
Since each step of Algorithm˜1 takes computation cost of $\mathcal{O}(NTd)$, this algorithm takes total computation cost of $\mathcal{O}(NTd)$ FLOPs. This implies diagonal SSD has the same computation cost as scalar-identity SSD (Dao and Gu, 2024, Theorem 3.8), and matches the maximum parameter complexity for SSD (Dao and Gu, 2024, Appendix A.1).
@@
The memory cost of the state data $A^{1},\cdots,A^{T},b_{1},\cdots,b_{T},c_{1},\cdots,c_{T}$ is $TN+TN+TN=\mathcal{O}(NT)$. In the first three steps of Algorithm˜1, each step generates $n$ matrices of size $T\times d$, and in the last step of Algorithm˜1, only one matrix of size $T\times d$ is generated. Therefore the memory cost of the intermediate step is $NTd+NTd+NTd+Td=\mathcal{O}(NTd)$. Considering all the memory costs above, we deduce that diagonal state-space duality has total memory cost $\mathcal{O}(NT)+\mathcal{O}(NTd)=\mathcal{O}(NTd)$.
@@
Note that the first three steps of Algorithm˜1 are operated in parallel for each $n\in[N]$, the diagonal state-space dual has separation into $N$ parallel computation processes, each of them costing time $\mathcal{O}(Td)$. Furthermore, from the definition of $f$ and $g$, note that each column of $X$ operates respectively during the whole processing of Algorithm˜1. Therefore the diagonal state-space dual has further separation into $Nd$ parallel computation processes, each process costing time $\mathcal{O}(T)$.
@@ 4.4 General SSMs with 1-SS Masked Attention Duals
The equivalence between $1$-$S$ and $1$-SSS-representable matrices appears in structured-matrix literature such as semiseparable, quasiseparable, and SSS forms (Vandebril et al., 2005). We extend it to $N$-$S$ matrices. Our proof is self-contained and constructive. It follows the causal setting and matches the attention mechanism in transformers. The result is more accessible to the ML community and supports our SSD design.
@@ 6 Conclusion
Structured state-space models (SSMs) (Dao and Gu, 2024; Gu and Dao, 2023) and attention mechanisms (Vaswani et al., 2017) represent two contrasting paradigms for sequence modeling. SSMs evolve a latent state step by step and run in linear time $O(T)$. Attention mixes all pairs and costs $O(T^{2})$. Structured State-Space Duality (SSD) (Dao and Gu, 2024) provides a theoretical bridge between these two views. We formalize and generalize the structured state-space duality between simple recurrent state-space models (SSMs) and masked attention mechanisms. We extend the duality introduced by (Dao and Gu, 2024) from the scalar-identity case to general diagonal SSMs (Section˜4.1) and show that these models retain the same computational complexity lower bounds while supporting richer dynamics (Section˜4.3). We further characterize the necessary and sufficient condition under which an SSM corresponds to a $1$-semiseparable masked attention mechanism (Section˜4.4). Lastly, we present two negative results (Section˜5): (i) this duality does not extend to standard softmax attention due to a rank explosion in the induced kernel; (ii) even for low-rank SSMs, such duality may still fail at representational level. Together, these results strengthen the theoretical bridge between recurrent SSMs and Transformer-style attention, and broaden the design space for expressive yet efficient sequence modeling.
@@ Acknowledgments
JH would like to thank Alex Reneau, Mimi Gallagher, Sara Sanchez, T.Y. Ball, Dino Feng and Andrew Chen for enlightening discussions on related topics, and the Red Maple Family for support.
JH is partially supported by Ensemble AI and Northwestern University. Han Liu is partially supported by NIH R01LM1372201, NSF AST-2421845, Simons Foundation MPS-AI-00010513, AbbVie, Dolby and Chan Zuckerberg Biohub Chicago Spoke Award.
@@ A.2 Condition for N-SS Matrix to Have Fine 1-SS Masked Attention Dual
for some $Q,K\in\mathbb{R}^{T\times N}$ and $L=\mathsf{1SS}(a_{1},a_{2},\cdots,a_{T})$ where $a_{1}a_{2}\cdots a_{T}\neq 0$.

### v2

# Last Update: December 11, 2025

@@
Last Update: December 11, 2025
@@
Jerry Yao-Chieh Hu†‡∗111jhu@u.northwestern.edu; jhu@ensemblecore.ai  Xiwen Zhang§∗222xiz284@ucsd.edu  Ali ElSheikh333alielsheikh2025@u.northwestern.edu  Weimin Wu444wwm@u.northwestern.edu  Han Liu†♯555hanliu@northwestern.edu
**footnotetext: These authors contributed equally to this work. Part of the work done during JH’s internship at Ensemble AI. Code is available at https://github.com/MAGICS-LAB/state_space_daulity.
@@ 1 Introduction
Nevertheless, while Dao and Gu (2024) suggest that analogues duality should hold for diagonal state-space models, a formal treatment in the SSM-attention setting appears to be missing. We provide a self-contained proof in in the SSM-attention language considered here. [^1] Similar equivalences are known in the structured-matrix (quasiseparable) literature, but they are not formulated in the SSM-attention language considered here. Building on it, we formalize diagonal structured state-space duality and give an explicit computation algorithm.
@@
  • General Diagonal SSMs (Sections˜4.1 and 4.2). We extend SSD beyond the simple $(\text{scalar})\times I_{N}$ state matrix to general diagonal state matrices. This enlarges the class of SSMs under the duality (from a single exponential decay to $N$ separate diagonal dynamics), thereby supporting richer sequence dynamics. [^2] For example, a diagonal SSM with two distinct decay rates can simultaneously capture both fast and slow time-scale patterns in a sequence, something a single-decay (scalar) SSM cannot do. In a toy regression task with a mixture of two exponential signals, a 2-dimensional SSM can fit both modes, whereas any 1-dimensional SSM fails (Section B.7).
@@
  • Higher-Rank Equivalence (Section˜4.4). We complement (Dao and Gu, 2024, Proposition 3.4) with a self-contained, constructive, strictly causal proof. It establishes the equivalence between (i) lower-triangular matrices of semiseparable rank at most $N$ and (ii) lower-triangular matrices admitting an $N$-dimensional sequential state-space representation, in the SSM-attention setting considered here. While the underlying equivalence is known in the structured-matrix (quasiseparable) literature, our formulation is tailored to causal SSM kernels and is reusable in subsequent duality results.
@@
We provide related work in Section˜2, and background in Section˜3. We show our main theory in Section˜4, and the limitations of structured state-space duality in Section˜5. Appendix˜B validates our theory with numerical studies.
@@ 2 Related Work
SSMs aim to model long-range dependencies with linear-time computation. S3 first introduced the idea of parameterizing state-space models with structured matrices to enable efficient sequence modeling (Gu et al., 2021). S4 introduces a state-space layer with stable diagonalization and fast convolution, which enabled long-context training and inference (Gu et al., 2021). S5 simplifies the design with a single multi-input multi-output SSM and preserved $O(T)$ scaling (Smith et al., 2023). Mamba adds input-dependent gating on the SSM projections and achieved strong accuracy with linear-time sequence modeling (Gu and Dao, 2023). These works establish SSMs as competitive sequence models and motivate analyses that compare their expressive power to attention.
@@ 4.2 Structured State-Space Duality for Diagonal SSMs with Full-Rank State Matrices
Remark 4.1 (An Example of “Richer Dynamics” of Diagonal SSM).
The analysis above shows substantial overlap between the kernels (equivalently, masked-attention matrices) representable by scalar-identity SSMs and by diagonal SSMs. However, diagonal SSMs are strictly more expressive: there exist kernels realizable by a diagonal SSM of state dimension $N$ that cannot be realized by any scalar-identity SSM with the same $N$. For example, let $N=2$, $T=4$, and consider
$M$ $=\begin{bmatrix}2&0&0&0\\ 1&2&0&0\\ 0&1&2&0\\ 0&0&1&2\end{bmatrix}=\begin{bmatrix}1&0&0&0\\ 1&1&0&0\\ 0&0&1&0\\ 0&0&1&1\end{bmatrix}+\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&1&1&0\\ 0&0&0&1\end{bmatrix}$.
This decomposition exhibits $M$ as a sum of two $1$-$S$ components, hence $M$ is realizable by a diagonal SSM with state dimension $N=2$ (equivalently, as a sum of two $1$-$S$ heads). In contrast, $M$ has $3>2=N$ new columns, so it cannot be realized by any scalar-identity SSM with state dimension $N=2$. In Section˜B.7, we further validate this “richer dynamics” phenomenon via multivariate time-series regression experiments.
@@ 4.3 Computational Complexity of Diagonal SSD
We define the functions $f:\mathbb{R}^{T}\times\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{T\times d}$ by $f(x,Y)_{:,s}=x\odot(Y_{:,s})$ for all $s\in[d]$, and $g:\mathbb{R}^{T}\times\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{T\times d}$ by $g(x,Y)_{1,:}=Y_{1,:}$ and $g(x,Y)_{t+1,:}=x_{t+1}\cdot g(x,Y)_{t,:}+Y_{t+1,:}$ for $t\in[T-1]$. Consider the SSM layer with state dimension $N$ defined by (3.2), where each $A^{t}$ is a diagonal matrix. The recurrence relation (3.2) also has representation
@@
  $Z^{n}\leftarrow f(b^{n},X)\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $O(NTd)$
  $H^{n}\leftarrow g(a^{n},Z^{n})\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $O(NTd)$
  $Y^{n}\leftarrow f(c^{n},H^{n})\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $O(NTd)$
  $Y\leftarrow\sum_{n=1}^{N}Y^{n}$ // Time $O(NTd)$
@@
Since each step of Algorithm˜1 takes computation cost of $O(NTd)$, this algorithm takes total computation cost of $O(NTd)$ FLOPs. This implies diagonal SSD has the same computation cost as scalar-identity SSD (Dao and Gu, 2024, Theorem 3.8), and matches the maximum parameter complexity for SSD (Dao and Gu, 2024, Appendix A.1).
Remark 4.2.
Specifically, we compute the $O(NTd)$ term as $NTd$ flops in each step and $4NTd$ flops in total. Therefore our algorithm not only has the same FLOP complexity order as SSD, but also exhibits a similar constant factor.
@@
The memory cost of the state data $A^{1},\cdots,A^{T},b_{1},\cdots,b_{T},c_{1},\cdots,c_{T}$ is $TN+TN+TN=O(NT)$. In the first three steps of Algorithm˜1, each step generates $n$ matrices of size $T\times d$, and in the last step of Algorithm˜1, only one matrix of size $T\times d$ is generated. Therefore the memory cost of the intermediate step is $NTd+NTd+NTd+Td=O(NTd)$. Considering all the memory costs above, we deduce that diagonal state-space duality has total memory cost $O(NT)+O(NTd)=O(NTd)$.
@@
Note that the first three steps of Algorithm˜1 are operated in parallel for each $n\in[N]$, the diagonal state-space dual has separation into $N$ parallel computation processes, each of them costing time $O(Td)$. Furthermore, from the definition of $f$ and $g$, note that each column of $X$ operates respectively during the whole processing of Algorithm˜1. Therefore the diagonal state-space dual has further separation into $Nd$ parallel computation processes, each process costing time $O(T)$.
Remark 4.3.
At present, there is no specialized kernel that directly accelerates diagonal-SSD computations. This is not a fundamental limitation: from a hardware perspective, diagonal SSDs admit the same asymptotic time complexity and parallelization structure as scalar-identity SSMs, and can in principle be implemented with similar accelerator support. Realizing such performance requires dedicated diagonal-SSD kernel design, which we leave to future work. The focus of this paper is to characterize the duality itself, rather than to develop optimized implementations.
@@ 4.4 General SSMs with 1-SS Masked Attention Duals
The equivalence between $1$-$S$ and $1$-SSS-representable matrices appears in structured-matrix literature such as semiseparable, quasiseparable, and SSS forms (Vandebril et al., 2005). We extend it to $N$-$S$ matrices. The equivalence between semiseparable matrices and structured state-space representations has been studied in the structured-matrix literature (e.g., under the notion of quasiseparable matrices). Our contribution here is not to claim a new mathematical equivalence, but to provide a self-contained, constructive, and strictly causal proof formulated entirely in the SSM/SSD notation. This formulation aligns directly with state-space models and masked attention, and is what enables us to use the equivalence as a building block in later results (e.g., Theorem˜4.1 and the diagonal SSM duality).
@@
Intuitively, a “new column” of $M$ is a column that is not in the span of the preceding columns. Equivalently, it introduces a new direction in the column space that cannot be synthesized from earlier columns. In the state-space view, each new column corresponds to an additional independent “memory direction” that must be carried by the latent state, i.e., information not representable by the previously accumulated state.
@@
Remark 4.7.
Mathematically, Lemma˜4.1 follows directly from Proposition˜4.2. Proposition˜4.2 establishes the structural form for the fine 1-SS case (Definition˜4.1). The general 1-SS case can be obtained by decomposing the matrix into several diagonal blocks, each of which is a fine 1-SS matrix. Applying Proposition˜4.2 to each block immediately yields Lemma˜4.1.
Remark 4.8 (Structural Nature of $1$-$S$ Duality).
Theorem˜4.1 is a structural theorem: it gives a maximally general necessary-and-sufficient condition for when an SSM admits a $1$-$S$ masked-attention dual. Since it applies to arbitrary SSMs (independent of parameterization, training objective, or optimization method) the statement is necessarily abstract and should not be read as a training-time recipe. Our proof is constructive in the mathematical sense: it identifies which subspaces (directions) must be preserved or can be merged to obtain a $1$-$S$ representation. The goal is to characterize the solution set, not to prescribe an implementation. Turning these structural insights into practical, hardware-efficient algorithms is an important direction for future work.
@@ 5 Limitations of Structured State-Space Duality
Remark 5.1.
Our rank-explosion obstruction is consistent with (Sieber et al., 2024). In their dynamical-systems formulation, an exact recurrent realization of softmax attention requires unbounded state dimension. In our terms, for generic $Q,K$, the matrix $\mathop{{\rm Softmax}}(QK^{\top})$ becomes (numerically) full rankfor large $T$ (not $N$-$S$ for any fixed $N$), so no finite-state SSM dual exists. Therefore, no finite-state SSM (equivalently, no fixed-size RNN) can represent softmax self-attention exactly.
@@ 6 Conclusion
Structured state-space models (SSMs) (Dao and Gu, 2024; Gu and Dao, 2023) and attention mechanisms (Vaswani et al., 2017) represent two contrasting paradigms for sequence modeling. SSMs evolve a latent state step by step and run in linear time $O(T)$. Attention mixes all pairs and costs $O(T^{2})$. Structured State-Space Duality (SSD) (Dao and Gu, 2024) provides a theoretical bridge between these two views. We formalize and generalize the structured state-space duality between simple recurrent state-space models (SSMs) and masked attention mechanisms.
Specifically, we extend the duality introduced by (Dao and Gu, 2024) from the scalar-identity case to general diagonal SSMs (Section˜4.1) and show that these models retain the same computational complexity lower bounds while supporting richer dynamics (Section˜4.3). We further characterize the necessary and sufficient condition under which an SSM corresponds to a $1$-semiseparable masked attention mechanism (Section˜4.4). We also present two negative results (Section˜5): (i) this duality does not extend to standard softmax attention due to a rank explosion in the induced kernel; (ii) even for low-rank SSMs, such duality may still fail at representational level. Lastly, we validate our results with numerical experiments (Appendix˜B). Together, these results strengthen the theoretical bridge between recurrent SSMs and Transformer-style attention, and broaden the design space for expressive yet efficient sequence modeling.
@@ Acknowledgments
JH thank Alex Reneau, Mimi Gallagher, Sara Sanchez, T.Y. Ball, Dino Feng and Andrew Chen for enlightening discussions on related topics, and the Red Maple Family for support. The authors would like to thank the anonymous reviewers and program chairs for constructive comments.
JH is partially supported by Ensemble AI, and the Walter P. Murphy Fellowship and the Terminal Year Fellowship (Paul K. Richter Memorial Award) of Northwestern University. Han Liu is partially supported by NIH R01LM1372201, NSF AST-2421845, Simons Foundation MPS-AI-00010513, AbbVie , Dolby and Chan Zuckerberg Biohub Chicago Spoke Award. This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding agencies.
@@ A.2 Condition for N-SS Matrix to Have Fine 1-SS Masked Attention Dual
for some $Q,K\in\mathbb{R}^{T\times N}$ and $L=\mathsf{1SS}(a_{1},a_{2},\cdots,a_{T})$ where $a_{1}a_{2}\cdots a_{T}\neq 0$, and suppose that $M$ has at least $N+1$ new columns.